OpenStudy (anonymous):

which would represent, cos theta(sec theta-cos theta), simplified as a single trig expression?

OpenStudy (anonymous):

Hi where are your options?

OpenStudy (anonymous):

1)cos^2 2)csc^2 3)sin^2 4)Sec^2

OpenStudy (anonymous):

\[\cos\theta(\sec \theta-\cos \theta) =\cos \theta \sec \theta-\cos^2 \theta\]is that right?

OpenStudy (anonymous):

thats not the equation but thats what it would equal when distributed

OpenStudy (anonymous):

do you know formula for the \(\sec \theta\) ?

OpenStudy (anonymous):

1/cos theta

OpenStudy (anonymous):

so\[\cos\theta(\sec \theta-\cos \theta) =\cos \theta \sec \theta-\cos^2 \theta\]\[=\cos \theta \frac{1}{\cos \theta}-\cos^2 \theta=1-\cos^2 \theta=\sin^2 \theta\]option 3 is your answer, is it clear for you?

OpenStudy (anonymous):

yes, thanks for the quick reply

OpenStudy (anonymous):

no problem :-)